Minutes of the Tracker Meeting

June 3, 2004


There was a discussion with Lowell and Ronaldo regarding increasing tower production capacity in order improve the schedule for delivery of the last flight tower.  Lowell remarked that this is a hot topic for project and mission office.  Since we are late getting started with flight production, we need to figure out where to squeeze to make commitments.  The proposal is for added capacity at G&A, adding tower assembly at G&A (in parallel with Pisa); and more fixtures at Plyform.   Quality still has to take precedence over schedule, but can any unessential tests be shortened or eliminated?  Preparation would have to start now to be ready in October with a second tower assembly line.  Plyform would be happy to go faster with the tray panel production, with more tooling.  1 extra INFN person will be needed to oversee testing at G&A. 


We need to understand better the schedule with and without the increased capacity.  Lowell suggested a meeting a week from tomorrow.


Ronaldo brought up an issue with sidewall prepreg procurement.  The sidewall costs have increased from original estimates of $200k to some $600k.  This causes a serious budget difficulty at INFN that SLAC may need to find a way to help with.  Ronaldo pointed out that the sidewalls originally were a SLAC responsibility but were traded for the Alcomp carbon-carbon material.  The carbon-carbon cost was in the neighborhood of $200k, so INFN got the bad end of the deal.


M4 screw material issue for side flexures: either SLAC must obtain the parts called out, or we need John Ku and Martin to review the specs for what can be purchased in Italy and then redline the drawing.

Pins for bottom tray closeout assembly: A286 steel is not available.  2mm pins arrive tomorrow in 304 stainless.  We need to know the annealing and other properties of this material.

Jim called for a 7:00am meeting Friday to go over these material issues.


Static test: Arthur reported that they are making good progress.  John is still doing some pre-analysis.  A preliminary test-readiness review will be held tomorrow after the MCM MRB.  Chris said that he would like to review the plan/procedure.  The official TRR will be Monday.  Jim will get a call-in number.


Another material issue is that we need to procure the correct fasteners for heat straps (the heat straps themselves still need to be procured).


The prepreg material is arriving at Plyform on Monday, enough for about 3 towers.

About $400k more is needed for prepreg for the remaining towers.  We are still waiting for a quotation from COI.


Mid tray panel status: 7 tray panels are completely finished through vibration (before painting).  8 more are ready to trim and paint.  Nusil will be in Plyform Monday.  15 panels have passed ESPI testing.  Production is expected next week for the remaining Tower-A trays.


TMCM pins have been obtained just now.  They are not plated yet.  INFN also received the non-flight MCM boards.  Sandro will be at G&A tomorrow and will know better the status after that.


Jerry and Albert will support getting bias circuits (78) and MCMs (29) ready to ship to Pisa with Peter A.


SLAC should buy the aluminum foil for sidewalls from Ohio and ship to Pisa asap.   Pisa will find out the oil content and other specs of their foil.


Jerry reported on the vibration fixture status: Tapemation quoted 8 to 10 weeks after receipt of material.  Humbolt quoted 6 weeks ARO and was much cheaper.  He suggests going to Humbolt for 1 asap.  Dave is working with the SLAC shop to get one built quicker. 


The flight flex circuit cable order was placed yesterday.  C0 cables are due in one week.


Studs are with Albert and should be getting coated today.  They will be done early next week, then send to LA for locking features.  Could be done by next Friday.


Nitronix hex nuts and floating nuts also need to be carried by Peter.  Jerry would like 8 of them for the vibe fixture as well.


Robert reported that 2 more EGSE systems are ready to ship to Italy.


I/F hardware drawings were released yesterday. 


T/V cables: Nicola said that Alenia needs 3 to 4 weeks to mate up the cables with the vacuum feedthroughs, so SLAC should hurry and send cables.